Use the five step process to determine each of the following areas


1. Between z = 0 and z = 2.67


2. Between z = -0.78 and z = -1.95


3. Between 1.27 and z = -2.1


4. To the right of z = 0.33


5. to the Left z = 1.96

1
Expert's answer
2022-03-01T09:20:33-0500

1.


"P(0<Z<2.67)=P(Z<2.67)-P(Z\le 0)"

"\approx0.9962074-0.5\approx0.496207"

2.


"P(-1.95<Z<-0.78)"

"=P(Z<-0.78)-P(Z\le -1.95)"

"\approx0.2176954-0.0255881\approx0.192107"

3.


"P(-2.1<Z<1.27)"

"=P(Z<1.27)-P(Z\le -2.1)"

"\approx0.8979577-0.0178644\approx0.880093"



4.


"P(Z>0.33)=1-P(Z\le0.33)"

"\approx0.370700"



5.


"P(Z<1.96)\approx0.975002"